With a 1500 MΩ resistor across its terminals, the terminal voltage of a certain battery is 2.60 V . With only a 5.11 Ω resistor across its terminals, the terminal voltage is 1.80 V

Part A

Find the internal emf of this battery.

Express your answer in volts to three significant figures.

Part B

Find the internal resistance of this battery.

Express your answer in ohms to three significant figures.

Part C

What would be the terminal voltage if the 5.11 Ω resistor were replaced by a 7.11 Ω resistor?

Express your answer in volts to three significant figures.
Answer #1

termical voltage = e - ir

and i = (e / r + R )


V = er + eR - er / (r+R)

V = eR / (r + R)

eR - Vr = VR


So,   1500e x 10^6 - 2.60r = 2.60 x 1500 x 106^6

and 5.11e - 1.80r = 5.11 x 1.80

solving


e = 2.27 volt ..Ans(A)

r = 2.6 ohm   ....Ans(B)


C) i = e / (r + R) = 2.27 / (7.11 + 2.6) = 0.234 A


V = e - ir = 2.27 - (2.6 x 0.234) = 1.66 volt